OpenStudy (anonymous):

Michelle rents a movie for a flat fee of $1.50 plus an additional $1.25 for each night she keeps the movie. If Michelle returned the movie and paid $4.00, how many nights did she keep the movie?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

one more: If g(x) = x^2 + 2, find g(3).

OpenStudy (anonymous):

\[1,5+1,25(x)-4=0\]

OpenStudy (anonymous):

\[g(3)=3^2+2\]

OpenStudy (anonymous):

1.25x + 1.5 = 4 1.25x = 4 - 1.5 1.25x =2.5 x = 2 nights.

OpenStudy (anonymous):

yes